# Binary-Report Peer Prediction for Real-Valued Signal Spaces

### v1

@@ 2.2 Equilibrium Characterization
In the case where the signal $x$ the agent receives satisfies $x\leq\tau$. Condition (2) simplifies to
@@ Results.
These thresholds exactly correspond to uninformative equilibria in the binary setting: if $\tau^{*}=\infty$, agents always report $L$, and if $\tau^{*}=-\infty$, agents always report $H$. Under some smoothness conditions on the predictive posterior function, we can furthermore characterize all non-infinite threshold equilibria. To do so, we define the function $G(x)=P(x;x)=\Pr[X^{\prime}\leq x\mid X=x]$ as the probability of another signal lying below $x$, conditioned on $x$. In Output Agreement, we find that threshold equilibria are characterized by thresholds $\tau$ such that $G(\tau)$ crosses 1/2. More precisely, we give the following necessary and sufficient conditions.
@@
Let finite threshold $\tau$ be given and $P(\tau;x)$ be continuous in $x$. If $\tau$ is a threshold equilibrium under the OA mechanism then $G(\tau)=1/2$. Conversely, if $G(\tau)=1/2$ and either (a) $P(\tau;x)$ is monotone decreasing in x or (b) $P(\tau;x)-1/2$ has a single crossing of 00 from positive to negative then $\tau$ is a threshold equilibrium under the OA mechanism.
@@
For intuition for the conditions, observe that the derivative of expected utility with respect to agent $i$’s choice of threshold $\tau$ in Equation (1) is $f(x)(2G(x)-1)$, where $f$ is the p.d.f. of $F$ and thus non-negative. Thus the necessary condition corresponds to the first order condition of optimizing the best response threshold and the sufficient conditions ensure this optimization is concave and quasiconcave respectively. However, the proof provided establishes optimality in the space of all possible strategies (which is required to be a threshold equilibrium), not merely that the chosen threshold is the optimal threshold.
@@ 2.3 Dynamics
We note that $\hat{\tau}(k\Delta t)$ is technically a best response to the mixture of thresholds at the previous time step $k-1$, rather than the probability distribution over thresholds. Ignoring this for the moment and taking the limit as $\Delta t\to 0$ of the discrete system in Equation 5, we end up with the following continuous best response dynamic:
@@
In most reasonable settings, one would expect $G(x)$ to limit toward 1 as $x$ approaches $\infty$, and toward 0 as $x$ approaches $\infty$- ∞. Specifically, as a signal $x$ becomes increasingly small, the probability another agent receives a signal smaller than $x$ should approach 0. Meanwhile, as a signal $x$ grows large, the probability another agent receives a signal smaller than $x$ should approach 1. As the quality of essay increases in a normally distributed class, for example, a grader would believe that the mass of essays below that quality should increase proportionally.
We show formally in § A that under such limit behavior of $G$, uninformative equilibria at $\tau^{*}=\pm\infty$ are stable in OA, while at least one non-infinite equilibrium exists which is unstable. In fact, one would expect in smooth, unimodal settings, as in § 2.4, that $G$ is monotone increasing. In this case, there are three equilibria, and by Theorem 2, the internal equilibrium is unstable while the uninformative equilibria are stable.
So what does this instability of internal equilibria mean for OA? When signals correspond monotonically to quality of an essay or task, agents will naturally move toward uninformative equilibria. These dynamics only further the fragility of truthfulness in OA: unless the threshold $\tau$ exactly balances out the conditional probabilities that other agents will report $H$ or $L$, agents will inevitably and increasingly misreport until they reach an uninformative consensus. There are multimodal Bayesian examples where several internal equilibria exist and some are thus topologically required to be locally stable (see § 6). However, as long as $G$ remains monotone increasing at extreme signals, we expect uninformative equilibria to be stable by our analysis in § A.
@@
Figure 2: Plot of $G-F$ in the Gaussian model (§ 2.4), for parameters $a=b=1$. $G(x)-F(x)$ is decreasing at $x=0$, meaning the corresponding equilibrium in DG is stable under the dynamics.
@@ 2.4 A Gaussian Model
[…]
@@
Existence of equilibria at $\tau=\pm\infty$ follows from Proposition 1. Note that $P(\tau;x)$ is strictly monotone decreasing and continuous in $x$ so that Theorem 1 applies. As $G(0)=\Phi(0)=\tfrac{1}{2}$, $\tau=0$ is an equilibrium for OA. Note also that $\Phi$ is strictly monotone, so $G(x)$ only crosses 1/2 once and $\tau=0$ is therefore the only non-infinite equilibrium (see a visualization in Figure 2). ∎
@@
[…]
@@
Let $c_{G}(\rho)=\frac{(1-\rho)}{b\sqrt{1+\rho}}$ be the coefficient of $x$ in $G(x)$. Then $G^{\prime}(x)=c_{G}(\rho)\phi(c_{G}x)$ for $\phi$ the PDF of the standard Normal, and so $G^{\prime}(0)=c_{G}(\rho)\phi(0)>0$. By Theorem 2, then, $\tau=0$ is unstable. Since 00 is the only non-infinite equilibrium, the equilibria $\tau=\pm\infty$ are stable by topological necessity. ∎
@@
3 Dasgupta-Ghosh
We next consider the mechanism proposed by Dasgupta and Ghosh [2013], which we will refer to as the Dasgupta-Ghosh (DG) mechanism. DG is a multi-task mechanism: agents participate in a set of peer prediction tasks, and their payment is a function of their reports across these tasks. For a given task, agent $i$ receives a payment of one if their report $r_{i}$ matches that of another agent $j$ on that task, as in Output Agreement. However, they are also penalized if their report matches that of the other agent on some unrelated task agent $i$ did not complete, with the intution that agents should be paid only for their performance over and above what would be expected by chance. [^3] There have been a variety of ways considered to calculate this penalty. This version corresponds to their $d=1$ option. In the binary signal setting, this penalty ensures that uninformative equilibria in OA where all agents agree to report H or L now have zero value, thereby ensuring strategic agents are not incentivized toward them.
@@ 3 Dasgupta-Ghosh
3.1 Equilibrium Characterization
For DG, the (ex-interim) expected utility for playing strategy $\sigma_{i}$ depends on $\sigma_{j}$ both directly in the bonus for agreement and indirectly through $\pi_{L}$:
The condition for a symmetric Bayes-Nash equilibrium can then be written as
For a threshold equilibrium, Condition 10 is equivalent to:
$\forall x\leq\tau,P(\tau;x)\geq\pi_{L}$ (11)
$\forall x>\tau,P(\tau;x)\leq\pi_{L}.$ (12)
Equilibrium results.
Now we have the tools to analyze threhsold equilibria in DG; we leave proofs to § B, since they largely follow the same logic as our proofs for OA. We first observe that uninformative strategies $\tau^{*}=\pm\infty$ remain equilibria in the DG mechanism.
Proposition 3.
$\tau^{*}=\pm\infty$ are both always threshold equilibria under DG.
Theorem 3.
Let finite threshold $\tau$ be given and $P(\tau;x)$ be continuous in $x$. If $\tau$ is a threshold equilibrium under the DG mechanism then $G(\tau)=F(\tau)$. Conversely, if $G(\tau)=F(\tau)$ and either (a) $P(\tau;x)$ is monotone decreasing in x or (b) $P(\tau;x)-\pi_{L}$ has a single crossing of 00 from positive to negative then $\tau$ is a threshold equilibrium under the DG mechanism.
3.2 Dynamics
We consider stability of equilibria for DG under the dynamics described by Equation (6). Again, we leave proofs in this section to § B. Note that as in OA, when everyone else is playing according to $\tau$, there exists a unique best response which is also a threshold strategy.
Proposition 4.
Now, as in OA, we can characterize the stability of equilibria.
Theorem 4.
@@ 3.2 Dynamics
The difference between $G$ in the OA case and $G-F$ here is important. We saw that $G$ was strictly increasing in natural settings like the Gaussian model. In contrast, in similar settings $G-F$ is strictly decreasing at a unique finite equilibrium point. While not all settings will behave as nicely as our Gaussian example, under reasonable behavior in the tails we expect existence of a stable nontrivial equilibria, while the uninformative equilibria remain unstable. We discuss such generalizations in § B.
3.3 Gaussian Model
Corollary 3.
[…]
Proof.
Uniqueness of $\tau=0$ as a finite threshold equilibrium follows. ∎
Now, consider the stability of these equilibria.
Corollary 4.
[…]
Proof.
As depicted in Figure 2, $G(x)-F(x)$ is strictly decreasing at 00. Formally, $G^{\prime}(x)-F^{\prime}(x)=c_{G}\phi(c_{G}x)-c_{F}\phi(c_{F}x)$, so $G^{\prime}(0)-F^{\prime}(0)=(c_{G}-c_{F})\phi(0)<0$ (since $c_{G}<c_{F}$ by (13)). Thus, by Theorem 4, the equilibrium at 00 is stable. Since $\tau=0$ is stable, the equilibria $\tau=\pm\infty$ are unstable by topological necessity. ∎
@@ 3.3 Gaussian Model
Since $\tau=0$ is the only stable equilibrium, it is in fact globally attracting: for any finite starting threshold $\tau(0)$, we have $\lim_{t\to\infty}\tau(t)=0$. In contrast to OA, this means that the behavior of DG in the Gaussian case is quite robust: unless the initial threshold starts exactly at an uninformative equilibrium it inevitably converges toward an informative one where $H$ and $L$ each get reported half the time. In this sense, DG behaves in a way one might hope based on its traditional analysis: it discourages uninformative equilibria.
@@
While robust, this equilibrium is also inflexible: the mechanism designer cannot specify a desired threshold, and must settle for $\tau=0$. Thus the semantics of $H$ and $L$ relative to the underlying signal are predetermined, as half of the tasks must be classified each way. In peer grading, for example, half the assignments would be labeled unsatisfactory, which might not be the desired outcome. [^4] See § 7 for futher discussion. We expect this general phenomenon, that there is a unique equilibrium threshold for DG which is determined by the underlying signal distribution, extends beyond the Gaussian case; see § 6.
4 Determinant-based Mutual Information (DMI) Mechanism
@@
The Determinant-based Mutual Information (DMI) mechanism was introduced by Kong [2024, 2020]. It is a minimal multi-task mechanism with several strong guarantees, such as being dominantly truthful under consistent strategies (see below). In the case of binary reports, the simplest version of the DMI mechanism collects 4 reports from a pair of agents, and computes a score based on a measure of their mutual information. To construct an unbiased estimator for this mutual information measure, the pairs of reports are split into two groups, and the determinants of the count matrices associated to each group are multiplied together.

### v2

@@ 2.2 Equilibrium Characterization
In the case where the signal $x$ the agent receives satisfies $x\leq\tau$, Condition (2) simplifies to
@@ Results.
These thresholds exactly correspond to uninformative equilibria in the binary setting: if $\tau^{*}=\infty$, agents always report $L$, and if $\tau^{*}=$-$\infty$, agents always report $H$. Under some smoothness conditions on the predictive posterior function, we can furthermore characterize all finite threshold equilibria. To do so, we define the function $G(x)=P(x;x)=\Pr[X^{\prime}\leq x\mid X=x]$ as the probability of another signal lying below $x$, conditioned on $x$. In Output Agreement, we find that threshold equilibria are characterized by thresholds $\tau$ such that $G(\tau)$ crosses 1/2. More precisely, we give the following necessary and sufficient conditions.
@@
Let finite threshold $\tau$ be given and $P(\tau;x)$ be continuous in $x$. If $\tau$ is a threshold equilibrium under the OA mechanism then $G(\tau)=1/2$. Conversely, if $G(\tau)=1/2$ and either (a) $P(\tau;x)$ is monotone decreasing in $x$ or (b) $P(\tau;x)-1/2$ has a single crossing of 0 from positive to negative then $\tau$ is a threshold equilibrium under the OA mechanism.
@@
The stronger monotonicity condition holds when a higher signal consistently increases the probability that a peer will receive a signal above the threshold. Meanwhile, the single-crossing condition allows for local belief fluctuations: a slightly higher signal might decrease confidence that a peer will report ‘H’, but there is no ambiguity regarding the optimal action. For example, in peer grading an agent may read an essay which is well-written but uses some amount of em-dashes: this may slightly decrease their confidence that someone else would give it a high mark because there is some chance it was generated by AI.
To better understand the conditions, we also observe that the derivative of expected utility with respect to agent $i$’s choice of threshold $\tau$ in Equation (1) is f(x)(2G(x)1)f(x)(2G(x)-1), where $f$ is the PDF of $F$ and therefore non-negative. Thus the necessary condition corresponds to the first order condition of optimizing the best response threshold and the sufficient conditions ensure this optimization is concave and quasiconcave respectively. However, the proof provided establishes optimality in the space of all possible strategies (which is required to be a threshold equilibrium), not merely that the chosen threshold is the optimal threshold.
@@ 2.3 Dynamics
We note that $\hat{\tau}(k\Delta t)$ is technically a best response to the mixture of thresholds at the previous time step k1k-1, rather than the probability distribution over thresholds. Ignoring this for the moment and taking the limit as $\Delta t\to 0$ of the discrete system in Equation 5, we end up with the following continuous best response dynamic:
@@
In most reasonable settings, one would expect $G(x)$ to limit toward 1 as $x$ approaches $\infty$, and toward 0 as $x$ approaches -$\infty$. Specifically, as a signal $x$ becomes increasingly small, the probability another agent receives a signal smaller than $x$ should approach 0. Meanwhile, as a signal $x$ grows large, the probability another agent receives a signal smaller than $x$ should approach 1. As the quality of essay increases in a normally distributed class, for example, a grader would believe that the mass of essays below that quality should increase proportionally.
We show formally in § A that under such limit behavior of $G$, uninformative equilibria at $\tau^{*}=\pm\infty$ are stable in OA, while at least one finite equilibrium exists which is unstable. In fact, one would expect in smooth, unimodal settings, as in § 2.4, that $G$ is monotone increasing. In this case, there are three equilibria, and by Theorem 2, the internal equilibrium is unstable while the uninformative equilibria are stable.
So what does this instability of internal equilibria mean for OA? When signals correspond monotonically to quality of an essay or task, agents will naturally move toward uninformative equilibria. These dynamics only further the fragility of truthfulness in OA: unless the threshold $\tau$ exactly balances out the conditional probabilities that other agents will report $H$ or $L$, agents will inevitably and increasingly misreport until they reach an uninformative consensus. There are multimodal Bayesian examples where several internal equilibria exist and some are thus topologically required to be locally stable (see § 4). However, as long as $G$ remains monotone increasing at extreme signals, we expect uninformative equilibria to be stable.
@@
Figure 2: Plot of GFG-F in the Gaussian model (§ 2.4), with $a=b=1$. G(x)F(x)G(x)-F(x) is decreasing at 0, meaning it a stable equilibrium.
@@ 2.4 A Gaussian Model
In the Gaussian model under OA, we have three equilibria at $\tau=0$ and $\pm\infty$.
@@
Existence of equilibria at $\tau=\pm\infty$ follows from Proposition 1. Note that $P(\tau;x)$ is strictly monotone decreasing and continuous in $x$ so that Theorem 1 applies. As $G(0)=\Phi(0)=\tfrac{1}{2}$, $\tau=0$ is an equilibrium for OA. Note also that $\Phi$ is strictly monotone, so $G(x)$ only crosses 1/2 once and $\tau=0$ is therefore the only finite equilibrium (see a visualization in Figure 2). ∎
@@
In the Gaussian model under OA, the threshold equilibrium $\tau=0$ is unstable, while the uninformative equilibria $\pm\infty$ are stable.
@@
Let $c_{G}(\rho)=\frac{(1-\rho)}{b\sqrt{1+\rho}}$ be the coefficient of $x$ in $G(x)$. Then $G^{\prime}(x)=c_{G}(\rho)\phi(c_{G}x)$ for $\phi$ the PDF of the standard Normal, and so $G^{\prime}(0)=c_{G}(\rho)\phi(0)>0$. By Theorem 2, then, $\tau=0$ is unstable. Since 0 is the only finite equilibrium, the equilibria $\tau=\pm\infty$ are stable by topological necessity. ∎
@@
3 Other Binary-Report Mechanisms
In this section, we characterize equilibria and their dynamics under the same real-valued signal model for three other popular binary-report peer prediction mechanisms.
3.1 Dasgupta-Ghosh
We consider the mechanism proposed by Dasgupta and Ghosh [2013], which we will refer to as the Dasgupta-Ghosh (DG) mechanism. DG is a multi-task mechanism: agents participate in a set of peer prediction tasks, and their payment is a function of their reports across these tasks. For a given task, agent $i$ receives a payment of one if their report $r_{i}$ matches that of another agent $j$ on that task, as in Output Agreement. However, they are also penalized if their report matches that of the other agent on some unrelated task agent $i$ did not complete, with the intution that agents should be paid only for their performance over and above what would be expected by chance. [^3] There have been a variety of ways considered to calculate this penalty. This version corresponds to their $d=1$ option. In the binary signal setting, this penalty ensures that the uninformative equilibria in OA where all agents agree to report H or L now have zero value, so that strategic agents are not incentivized toward them.
Model.
@@ 3.1 Dasgupta-Ghosh
Equilibrium characterization and dynamics.
We leave formal derivations of equilibrium characterization and dynamics to § B, because they follow in a similar way to our OA analysis. Importantly, while the uninformative equilibria still exist, the penalty term induces a different characterization of finite equilibria: under the same continuous and single-crossing assumptions over the predictive posterior, $\tau^{*}$ is an equilibrium if and only if $G(\tau^{*})=F(\tau^{*})$ (where as a reminder, $F(\tau^{*})=\Pr[X\leq\tau^{*}])$. Moreover, $\tau^{*}$ is stable if and only if $G(\tau)-F(\tau)$ is strictly increasing at $\tau^{*}$.
@@
The difference between $G$ in the OA case and GFG-F here is important. We saw that $G$ was strictly increasing in natural settings like the Gaussian model. In contrast, in similar settings GFG-F is strictly decreasing at a unique finite equilibrium point. We find under the Gaussian model, in fact, that the same equilibria exist at $(0,\pm\infty)$; but 0 is stable, while the uninformative equilibria are unstable (see Figures 2 and 2). Since $\tau=0$ is the only stable equilibrium, it is in fact globally attracting: for any finite starting threshold $\tau(0)$, we have $\lim_{t\to\infty}\tau(t)=0$.
@@
In contrast to OA, this means that the behavior of DG in the Gaussian case is quite robust: unless the initial threshold starts exactly at an uninformative equilibrium it inevitably converges toward an informative one where $H$ and $L$ each get reported half the time. In this sense, DG behaves in a way one might hope based on its traditional analysis: it discourages uninformative equilibria. While not all settings will behave as nicely as our Gaussian example, under reasonable behavior in the tails we expect existence of a stable nontrivial equilibria, while the uninformative equilibria remain unstable. We discuss such generalizations in § B.
@@
While robust, the stable equilibrium in DG under the Gaussian model is also inflexible: the mechanism designer cannot specify a desired threshold, and must settle for $\tau=0$. Thus the semantics of $H$ and $L$ relative to the underlying signal are predetermined, as half of the tasks must be classified each way. In peer grading, for example, half the assignments would be labeled unsatisfactory, which might not be the desired outcome. [^4] See § 6 for futher discussion. We expect this general phenomenon, that there is a unique equilibrium threshold for DG which is determined by the underlying signal distribution, extends beyond the Gaussian case; see § 4.
3.2 Robust Bayesian Truth Serum
The Robust Bayesian Truth Serum (RBTS) mechanism [Witkowski and Parkes, 2012], inspired by the Bayesian Truth Serum (BTS) [Prelec, 2004], is a non-minimal mechanism: each agent $i$ submits both an information report $r_{i}\in\{L,H\}$ and also a prediction report $p_{i}\in[0,1]$. The prediction report represents agent $i$’s belief about the frequency of $H$ reports from all agents. The mechanism incentivizes prediction reports using the Brier score $S(p,r)=2p\mathbf{1}[r=H]-p^{2}$ (also known as the quadratic score), which is an example of a strictly proper scoring rule [Brier, 1950, Gneiting and Raftery, 2007]. RBTS randomly picks a reference agent $j$ and peer agent $k$, and pays agent $i$
where t(L)=1t(L)=-1 and $t(H)=1$, and $\delta>0$ is chosen by the mechanism.
We leave our formal analysis of RBTS to § C, where we (perhaps surprisingly) are able to characterize equilibria under similar monotonicity and continuity conditions as OA and DG. We find equilibria coincide with all thresholds that satisfy $Q(\tau)=G(\tau)$, where $Q(x)=\mathop{\mathbb{E}}_{x^{\prime}\sim\beta(x)}P(x;x^{\prime})$. That is, $Q$ represents the expected posterior of the agent’s peer at signal $x$, conditional on $x$. We then show that the location and dynamics of equilibria in RBTS under real-valued signals match DG for the Gaussian model (though convergence to stable equilibria is faster under RBTS), meaning the same inflexibility is inherited.
3.3 Determinant-based Mutual Information (DMI) Mechanism
@@
The Determinant-based Mutual Information (DMI) mechanism was introduced by Kong [2024, 2020]. It is a minimal multi-task mechanism with several strong guarantees, such as being dominantly truthful under consistent strategies. In the case of binary reports, the simplest version of the DMI mechanism collects 4 reports from a pair of agents, and computes a score based on a measure of their mutual information. To construct an unbiased estimator for this mutual information measure, the pairs of reports are split into two groups, and the determinants of the count matrices associated to each group are multiplied together.